Colony Inspection 28th May

Observations:

weather back to normal for time of year – temperatures up to 18C.

 

flowering: chesnut finishing, wisteria, thyme and elder just starting

 

Blue – new colony from Wilson/Pagden has built QCs on frame from Aqua after colony was re-arranged and old queen moved to nuc. Also Denim has started swarming with 4 charged QCs, so took all down and will introduce a sealed QC from Blue.

2nd June – Aqua – cut out a sealed QC and put into Denim. Moved Denim queen to nuc and put in River Mole site, added sealed QC plus 3 frames foundation to replace frames added to nuc. Took down all sealed and unsealed QCs – also removed DT and found 50 in 200 mites – so will need to consider treating. But may get away with it as there will be a period of no new brood because of re-queening activity.

Hive Record
Colony
Q
BIAS
Brood
Stores
CommentNext Inspection
AquaYY2 BIAS including eggs,OK 1 egg in queen cupLeave alone for now
BlueNY3 BIAS, sealed.6 sealed QCs and 4 unsealed - took down to 2About 2 KGs in superRe-arranged into 1 brood boxTake 1 QC to Denim
CyanYY8 + 7About 4Kg in super, going wellStarting to lay up DT. Re-arranged brood box to give more spaceRemove DT and add 2nd super
DenimYY8 BIAS, 4 charged QCs2 Kgs in superSwarming, so took down all QCs and will add sealed cell from BlueRemove queen to nuc
EmeraldYY7 + 8 About 2 Kgs in super2 in 150 in DT - varroa under control, remarked queenCheck when 2nd super needed
FlaxYY6 BIAS not expandingOK but not working super25 in 200 in DT - need to manage VarroaMay be candidate for unite to one of the nucs
GreenYY6 BIAS, so has started to expandOKNice brood patternCheck when super is required
HoneydewYY7 BIAS2 Kgs in superQueen not marked or clipped ! Lots of drone brood scattered through the colonyCheck when next super required.